• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Feuerbach

    Log entry time 13:35:06 on July25,2003

    Entry number 106845

    keyword=S2m ADC signal investigations

    The ADC signals for Left HRS S2-9R and 15L were checked. The signals came back after disconnecting/reconnecting at the splitter box (9R) and lower patch panel (15L). These ADC signals were lost after replacing the ADC cables yesterday. This did not affect the TDC or trigger signals, which have remained intact.

    At this point, a short cosmic run shows that the ADC and TDC signals from all channels are good.